A new packing machine will be purchased at a cost of $57,000. The old machine that is being replaced is going to be sold today for $5000. The salvage of the new machine is $7500 after its 10 year useful life. If the new machine reduces annual expenses by $5000, what is the present worth of the new machine, assuming 25% return on investment.
Answer #1

MARR = 25%

net initial cost = 57000 - 5000 = 52000

Annual savings = 5000

salvage = 7500

t = 10 yrs

PW = -52000 + 5000*(P/A,25%,10) + 7500*(P/F,25%,10)

= -52000 + 5000*3.570503 + 7500*0.107374

= -33342.17